# Discovering the underlying analytic structure within Standard Model constants using artificial intelligence

### v2

@@ Abstract
This paper presents a search for underlying analytic structures among the fundamental parameters of the Standard Model (SM) using symbolic regression and genetic programming. We identify the simplest analytic relationships connecting pairs of these constants and report several notable expressions obtained with relative precision better than 1%. These results may serve as valuable inputs for model builders and artificial intelligence methods aimed at uncovering hidden patterns among the SM constants, or potentially used as building blocks for a deeper underlying law that connects all parameters of the SM through a small set of fundamental constants.
@@ 1 Introduction
The Standard Model (SM) remains highly successful in describing fundamental particles and their interactions, accurately predicting a broad range of experimental results. However, it is still considered a model rather than a complete theory, as it relies on approximately 20 fundamental parameters that must be determined experimentally [1]. This reliance limits its predictive power. One of the most important open questions in the SM is why these constants have the specific values we observe, and whether a deeper theoretical foundation might explain them.
@@
The Large Hadron Collider (LHC) has enabled numerous stringent tests of the SM, yet no evidence of new physics has been observed. Most searches for new physics at the LHC rely heavily on the SM’s predictive power. However, the SM contains numerous adjustable parameters that must be determined experimentally, which limits the interpretability of the measured cross sections. Moreover, the SM offers limited predictive power for many categories of events produced in $pp$ collisions, leaving significant regions of event topology still largely unexplored [2].
@@
Attempts to identify relationships among the parameters of the SM have a long history [3, 4, 5]. However, no convincing evidence has been found to support the existence of such relationships. As a result, these constants are generally regarded as free parameters of the SM, with no apparent underlying connections.
Finding analytic relationships through brute-force searches or by noticing coincidences in numerical values or measurements was characteristic of the early stages of science - before the development of fundamental theories that could explain connections between seemingly unrelated quantities. However, this empirical approach proved to be useful and sometimes led to the discovery of new laws that revealed the underlying reasons for such coincidences. It is not improbable that a similar strategy could be valuable today, especially with the increasing availability of artificial intelligence (AI) algorithms and greater computational power.
@@
In this paper, we explore the possibility of uncovering underlying structures among the fundamental constants of the SM using symbolic regression and genetic programming - a type of evolutionary algorithm inspired by natural selection and rooted in the broader field of AI. We apply this method to investigate whether analytic relationships among SM constants reveal an inherent structure or potentially point to a deeper, unifying law that governs them.
@@ 2 Technical details
Genetic programming (GP) is a powerful technique in symbolic regression that evolves analytic expressions directly from data [6]. By representing mathematical expressions as trees and applying evolutionary algorithms, GP explores a wide solution space unconstrained by predefined model structures, enabling the discovery of complex, previously hidden relationships.
In symbolic regression, the system learns an (symbolic) expression from a set of input-output pairs, often many such pairs, and often there are many input values (representing the dependent variables), and a single output (the independent variable). The generated expressions then contain relations with these variables, functions of these variables, as well as constants. It can then be used to evaluate expressions with inputs not seen in the given data set.
The symbolic computations were performed using a program written in the Picat language [7], a logic-based, multi-paradigm programming language that combines features from logic programming, constraint solving, functional programming, planning, and imperative programming.
The program primarily leverages Picat’s logic programming (non-determinism), functional constructs, and imperative features. It is available on the second author’s website [8].
The specific variant of symbolic regression that is used here (constant identification), is different in some important aspects:
  • The data set for a single run/experiment consists of a single input-output pair:
    <1, value of the physical constant>
    where 1 is the input, and the value of the specific physical constant is the output.
  • The algorithm then tries to learn some relation between the constant 1 and value the physical constant, generating an expression. See below for more on the specific genetic programming method.
  • The found expression does not contain variable names as described above for the general symbolic regression approach. In our version of constant identification, everything in the generated expression is either a function or constants (physical constant or an integer).
  • The equation for the physical constants that is shown below in the paper contains the name of the physical constants, but that is handled by post-processing after the symbolic regression part is done.
@@
    • functions: +, --, *, /, $\sqrt{(}x)$, $x^{2}$, $x^{3}$, $x^{4}$, $\exp(x)$, $\log(x)$;
    • the integers from 1 to 10;
@@
    The use of only small integers instead of larger integers or random floats is for simplicity and for the interpretability of an expression.
  • The analytic formulas are evaluated for each expression in a population, and the results that are closest to the output value are kept for the next generation. The formulas that are too far are replaced with previous expressions that are combined from other expressions using the standard genetic programming operators crossover and mutation (with some specific probability). New random formulas may also be added (from the same set as above).
  • When an expression is sufficiently close to the value of output $c$ within the allowed error of the output - then this is considered a solution. All solutions found within the allotted number of runs (1000) are reported. If no solution is found within the allotted number of runs, the program continues to run until a solution is found (and then reported), or until the program times out. The timeout for each experiment, based on a given output–input configuration, was set to 60 seconds.
@@
The results presented in this paper are based on 3,249 distinct runs, each involving the testing of an output against one or two input constants, as will be discussed later. Many single runs produced multiple solutions. The results were obtained from multiple runs with different random seed values, which takes into account the inherent stochastic nature of GP. Duplicate solutions were removed during the post-processing step. All runs took in total about 15,000 CPU hours with the Intel E5-2650 v4 processor and 252Gb RAM.
@@ 2.1 Inputs for the genetic algorithm
A critical aspect of evaluating analytic relationships between SM parameters is the selection of an appropriate input space for the algorithms. Working with SM parameters presents several challenges. First, most of these parameters have physical units, which restricts the applicability of the algorithms due to dimensional analysis constraints — only certain analytic structures are valid when dimensions are taken into account. To enable numerical evaluation, we focus on dimensionless constants. There are arguments [9] suggesting that the laws of physics should be independent of the choice of units or measuring instruments. Dimensionless constants, in particular, are well-suited for probing the deep structure underlying physical laws.
The second challenge for genetic algorithms lies in the wide range of SM values, spanning several orders of magnitude — from the electron mass to the top quark mass, the latter being the heaviest among all SM particles.
Finally, the uncertainties associated with the SM parameters vary significantly and must be taken into account when evaluating the robustness of any discovered relationships. If one variable has very small uncertainties, it offers little advantage if other input variables have large uncertainties, which will dominate the overall error.
@@
One way to address the issue of physical units in the SM is to normalize all particle masses by dividing them by another parameter that has units of mass. However, this approach introduces a degree of arbitrariness in the choice of the reference mass. A possible option for the divider is the Planck mass [1], $m_{\mathcal{P}}=1.22089(6)\times 10^{19}$ GeV. While theoretically appealing, this choice is impractical for our purposes, as the GP algorithm we employ cannot handle values smaller than $10^{-16}$.
@@
To resolve this, we normalize all SM masses by the mass of the $\rho(770)$ meson, 0.77526±0.000230.77526$\pm$ 0.000230.77526 ± 0.00023 GeV [1]. At first glance [^2] We could also use the mass of any fundamental particle for normalization, such as the electron mass. However, that particle would then be excluded from our GP analysis., this might seem like an unconventional choice. However, it is not without justification for the GP algorithm. The $\rho(770)$ meson holds a special place in the context of quantum chromodynamics, plays a key role in hadronic structure, and even electromagnetic interactions. Its mass lies conveniently in the GeV range, which allows us to rescale the SM mass spectrum into a range well-suited for GP.
@@
For example, the mass of the electron, 0.51099895000(15) MeV, has an exceptionally small relative uncertainty of $3\times 10^{-8}\%$, while the Higgs boson mass, 125.20(11)125.20(11)125.20 ( 11 ) GeV [1], has a much larger relative uncertainty of approximately 0.09%. This disparity in precision poses a challenge for numerical algorithms. However, scaling all masses by the $\rho(770)$ meson mass and propagating its uncertainty to the resulting ratios helps mitigate this issue.
Thus, in our GP algorithm, the maximum precision of the input parameters was set to 0.03%. The input includes two dimensionless constants whose precision significantly exceeds this threshold: the mathematical constant $\pi$, and the fine-structure constant $\alpha$ (or its inverse value $\alpha^{-1}$, which will be used in our study). Although the constant $\pi$ is not a parameter of the SM, it frequently appears in the laws of physics. To maintain consistency, we reduce their precision to $\pi=3.14159(1)$ and $\alpha^{-1}=137.036(1)$. Even with this adjustment, they remain 40–100 times more precise than the most precisely known mass parameter — the electron mass normalized by the $\rho(770)$ meson mass. Increasing the precision of $\pi$ and $\alpha^{-1}$ is not expected to affect the results involving particle masses, as the overall uncertainty will still be dominated by uncertainties in the mass values. However, using overly precise constants can introduce numerical instability and make it more difficult for GP to reproduce known relationships. These validation tests will be discussed later.
With all these assumptions, the input parameters for the GP processing are shown in Table 1. Note that we did not include the PMNS mixing angles [1], which are often considered a part of the fundamental constants of SM. They are more relevant for neutrino physics, but neutrino masses are not considered. The measured precision of their values is rather low which can trigger many “noisy” solutions in the evaluation of relations using GP. We also did not include the speed of light ($c$), which has units. However, one should mention that $c$ is indirectly included in the calculation via the fine-structure constant $\alpha$.
Table 1: Dimensionless constants [1] used for the GP algorithm. All particle masses in this table are divided by the mass 0.77526±0.000230.77526$\pm$ 0.000230.77526 ± 0.00023 GeV of the $\rho(770)$ meson to improve the value range and for reducing the large variability of uncertainties, making them more uniform for GP. The traditional symbol $m$ was chosen for the rescaled masses to simplify the presentation of the results. We require that the uncertainties for $\pi$ and $\alpha^{-1}$ are at most a factor 100 higher than for best measured mass. Making them more precise should not contribute to the results. The table presents both the absolute measured uncertainties (±$\pm\varepsilon$) and the relative uncertainties (±$(\pm\varepsilon^{rel}$, in percent).
Constant Name Value ±$\pm\varepsilon$$\pm\varepsilon^{rel}$ (%)
PI $\pi$ 3.14159 1e-05 0.0003
Fine-structure constant (inv) $\alpha^{-1}$ 137.036 0.001 0.0007
Strong coupling constant at $Z^{0}$ $\alpha_{S}$ 0.1180 0.0009 0.7627
CKM 12-mixing angle $\theta_{12}$ 0.22501 0.00068 0.3022
CKM 23-mixing angle $\theta_{23}$ 0.04183 0.00079 1.8886
CKM 13-mixing angle $\theta_{13}$ 0.003732 9e-05 2.4116
CKM CP-violating Phase $\delta$ 1.147 0.026 2.2668
electron mass $m_{e}$ 0.0006591 2e-07 0.0303
muon mass $m_{\mu}$ 0.1363 4e-05 0.0293
tau mass $m_{\tau}$ 2.292 0.0007 0.0305
u-quark mass $m_{u}$ 0.002786 9e-05 3.2304
d-quark mass $m_{d}$ 0.006062 9e-05 1.4847
s-quark mass $m_{s}$ 0.1206 0.001 0.8292
c-quark mass $m_{c}$ 1.642 0.0059 0.3593
@@
The least precise masses among our inputs are those of the light-flavor quarks, which were taken from the PDG using the $\overline{\text{MS}}$ scheme at a renormalization scale of 2 GeV. The most significant challenge is also when using the CKM mixing angles as inputs. Due to their relatively large uncertainties, there is a considerable risk of generating spurious equations.
It should be emphasized that the presence of the mass of $\rho(770)$ (or any other particle mass used for obtaining dimensionless SM constants) should be considered a handy method for deriving short relationships in GP. It is an auxiliary parameter. The final theory can be free of this constant after successful variable substitutions, which should be performed after obtaining the analytic expressions.
2.2 Search strategy
The input parameters shown in Table 1 will be used for the following studies as described below.

### v3

@@ Abstract
This paper presents a method for uncovering hidden analytic relationships among the fundamental parameters of the Standard Model (SM), a foundational theory in physics that describes the fundamental particles and their interactions, using symbolic regression and genetic programming. Using this approach, we identify the simplest analytic relationships connecting pairs of these constants and report several notable expressions obtained with relative precision better than 1%. These results may serve as valuable inputs for model builders and artificial intelligence methods aimed at uncovering hidden patterns among the SM constants, or potentially used as building blocks for a deeper underlying law that connects all parameters of the SM through a small set of fundamental constants.
@@ 1 Introduction
The Standard Model (SM) is the theory in physics that describes the fundamental particles and their interactions. It remains highly successful in predicting a wide range of experimental results. However, it is still considered a model rather than a complete theory, as it relies on approximately 20 fundamental parameters that must be determined experimentally [ParticleDataGroup:2024cfk]. For example, the masses of the most fundamental particles—such as the six quarks (denoted by the letters $u$, $d$, $s$, $c$, $b$, and $t$), the vector bosons ($Z$ and $W$), the Higgs boson ($H$), and the leptons (electrons $e$, muons $\mu$, and tau leptons $\tau$)—are treated as free parameters that must be determined experimentally. This reliance on such parameters limits the predictive power of the SM. One of the most important open questions in the SM is why these constants have the specific values we observe, and whether a deeper theoretical framework could explain them.
@@
The Large Hadron Collider (LHC) has enabled numerous stringent tests of the SM, yet no evidence of new physics has been observed. Most searches for new physics at the LHC rely heavily on the SM’s predictive power. However, the SM contains numerous adjustable parameters that must be determined experimentally, which limits the interpretability of the measured cross sections. Moreover, the SM offers limited predictive power for many categories of events produced in $pp$ collisions, leaving significant regions of event topology still largely unexplored [universe10110414].
@@
Attempts to identify relationships among the parameters of the SM have a long history [WBook, Nielsen:1994ab, Froggatt__2003]. However, no convincing evidence has been found to support the existence of such relationships. As a result, these constants are generally regarded as free parameters of the SM, with no apparent underlying connections.
When there is no available theory that can describe apparently unconnected constants, the first step is to identify the simplest analytic relationships between their values, before constructing a theory based on specific physics principles. Finding analytic relationships through brute-force searches by noticing coincidences in measured values is common for science—even before the development of fundamental theories that could explain connections between seemingly unrelated quantities. This empirical approach proved to be useful and sometimes led to discoveries of new laws that revealed the underlying reasons for such coincidences. For example, in 1885, Johann Balmer derived an empirical formula that accurately described the wavelengths of the visible spectral lines of hydrogen. At the time, it was regarded solely as a mathematical relation, lacking a corresponding physical interpretation. But this mathematical regularity was later explained by quantum mechanics through Bohr’s model of the atom in 1913.
A similar strategy may be valuable for studying the many free constants of the SM. Compared to the earlier example, the main challenges lie in the large number of parameters, the presence of physical units for some of them, and the wide variation in measurement precision, since some constants are known with extremely high accuracy, while others are not. Today, however, the search for numerical regularities in many seemingly unrelated parameters is becoming increasingly feasible thanks to advances in artificial intelligence (AI) and growing computational power.
@@
In this paper, we present an algorithm designed to uncover potential underlying structures among the fundamental constants of the SM using symbolic regression and genetic programming (GP)—a type of evolutionary algorithm inspired by natural selection and rooted in the broader field of AI. We apply this method to explore whether analytic relationships among SM constants may hint at an intrinsic structure or possibly point toward a unifying law that governs them. We also discuss how our method may be adapted for application in a broad range of scientific fields.
2 The Method
Our AI-based approach to addressing the problem of free parameters in the SM proceeds as follows:
  1. 1.
    The first step is a traditional symbolic regression [koza92], which derives analytical expressions from SM parameter data, generating a large set of candidate relations. This step is the most computationally demanding and produces large datasets containing analytical relationships among SM constants. Such datasets inevitably include spurious analytic expressions that arise purely from numerical coincidences rather than genuine physical connections.
  2. 2.
    The second step operates on the datasets generated in the first stage. At this stage, dimensional analysis and general expectations from the SM are applied to identify the most plausible analytical structures within a high-dimensional functional space. We expect that some constraints can be applied to filter out numerical noise and retain only relations that satisfy certain principles. The obtained relations may provide hints of symmetries or dynamical models underlying the SM. However, since dimensional analysis has not yet been implemented in the GP algorithm, this step cannot currently be performed in a straightforward manner.
This paper describes the first step in full detail, i.e., the creation of datasets of analytic relationships between the SM parameters generated by the GP algorithm. The purpose of this step is to provide analytical relationships among fundamental physical constants, some of which can be used as building blocks for models with a reduced number of free parameters. These datasets can also serve as input for various AI techniques to analyze patterns and structures that may reveal hidden relationships among SM parameters. We will also provide an example of the second step based on some basic SM expectations. However, a full analysis involving this step should be presented in a dedicated physics analysis paper, which is beyond the scope of this article.
@@ 3 Technical details
Genetic programming is a powerful technique in symbolic regression that evolves analytic expressions directly from data [koza92]. By representing mathematical expressions as trees and applying evolutionary algorithms, GP explores a wide solution space unconstrained by predefined model structures, enabling the discovery of complex, previously hidden relationships.
In symbolic regression, the system learns an (symbolic) expression from a set of input–output pairs, often many such pairs, and often there are many input values (representing the dependent variables), and a single output (the independent variable). The generated expressions then contain relations with these variables, functions of these variables, as well as constants. It can then be used to evaluate expressions with inputs not seen in the given data set.
The symbolic computations were performed using a program written in the Picat language [PicatWebsite], a logic-based, multi-paradigm programming language that combines features from logic programming, constraint solving, functional programming, planning, and imperative programming. The program primarily leverages Picat’s logic programming (non-determinism), functional constructs, and imperative features. It is available on the second author’s website [PicatSymbolicRegression].
The specific variant of symbolic regression used here is designed for constant identification. This method is different from the traditional symbolic regression in some important aspects:
  • The data set for a single run/experiment consists of a single input-output pair: $(a,b)$, where $a$ is an input value, and $b$ is a value of the expected output.
  • The algorithm then tries to learn some relation between the constant $a$ and $b$, generating an analytic expression. See below for more on the specific genetic programming method.
  • The found expression does not contain variable names as described above for the general symbolic regression approach. In our version of constant identification, everything in the generated expression is either a function or constants (a physical constant or an integer).
  • The equation for the physical constants that is shown below in the paper contains the name of the physical constants, but that is handled by post-processing after the symbolic regression part is completed.
@@
    • functions: +, -, *, /, $\sqrt{x}$, $x^{2}$, $x^{3}$, $x^{4}$, $\exp(x)$, $\log(x)$, $\cos(x)$, $\sin(x)$, $\tan(x)$, $\arctan(x)$;
    • the integers from 1 to 10. The use of only small integers instead of larger integers or random floats is for simplicity and for the interpretability of an expression;
@@
  • The analytic formulas are evaluated for each expression in a population, and the results that are closest to the output value are kept for the next generation. The formulas that are too far from the output value are replaced with previous expressions that are combined from other expressions using the standard genetic programming operators crossover and mutation (with some specific probability). New random formulas may also be added (from the same set as shown above).
  • When an expression is sufficiently close to the value of output cc—within the allowed error of the output—then this is considered a solution. All solutions found within the allotted number of runs (1000) are reported. If no solution is found within the allotted number of runs, the program continues to run until a solution is found (and then reported), or until the program times out. The timeout for each experiment, based on a given output–input configuration, was set to 60 s.
@@
The results presented in this paper are based on 3249 distinct runs using inputs either with or without mass units. Each involving the testing of an output against one or two input constants, as will be discussed later. Many single runs produced multiple solutions. The results were obtained from multiple runs with different random seed values, which takes into account the inherent stochastic nature of GP. Duplicate solutions were removed during the post-processing step. All runs took, in total, about 40,000 CPU hours with the Intel E5-2650 v4 processor and 252 Gb RAM.
@@ 3.1 Inputs for the genetic algorithm
A critical aspect of evaluating analytic relationships between SM parameters is the selection of an appropriate input space for the algorithms.
@@
Working with SM parameters presents several challenges. First, most of these parameters have physical units, which restricts the applicability of the algorithms due to dimensional analysis constraints—only certain analytic structures are valid when dimensions are taken into account.
3.2 Inputs with Units
Table 1 lists the fundamental SM parameters commonly used in particle physics, as reported by the PDG [ParticleDataGroup:2024cfk]. These constants, including the masses of fundamental particles expressed in physical units (e.g., MeV), are used as inputs for GP. We will discuss these constants below.
We did not include the PMNS mixing angles [ParticleDataGroup:2024cfk] in the GP algorithm. These parameters are regarded as part of the fundamental constants of the SM. They are primarily relevant to neutrino physics, but neutrino masses are not considered in this paper since, at present, only upper limits on their values exist. In addition, the experimental precision of the PMNS mixing angles remains relatively low [Esteban:2024eli], with uncertainties ranging from 5% to 30%, which is significantly worse than for the other SM constants. Including such poorly determined quantities would generate a large number of “noisy” analytic solutions in the GP analysis, making them difficult to evaluate within the limited CPU resources available for this study, and potentially obscuring more meaningful relations. Nevertheless, the PMNS mixing angles may be incorporated in future work when greater computational resources become available.
We also did not include cosmological constants and the speed of light ($c$). However, one should mention that $c$ is indirectly included in the calculation via the fine-structure constant $\alpha$.
Table 1: Physics fundamental constants from the PDG [ParticleDataGroup:2024cfk] used for the GP algorithm. We require that the uncertainties for $m_{e}$, $\pi$ and $\alpha^{-1}$ are at most a factor 100 higher than for the Higgs boson mass. Making them more precise should not contribute to the results. The second half of the table shows the masses of the fundamental particles in MeV units. The table presents both the absolute uncertainties ($\pm\varepsilon$) and the relative uncertainties $(\pm\varepsilon^{rel}$, in percent). When the PDG reports an asymmetric uncertainty, the largest value is used.
Constant Name Value $\pm\varepsilon$ $\pm\varepsilon^{rel}$ (%)
PI $\pi$ 3.14159 $10^{-5}$ 0.0003
Fine-struct. (inv) $\alpha^{-1}$ 137.036 0.001 0.0007
$\alpha_{s}$ at $Z^{0}$ $\alpha_{S}$ 0.1180 0.0009 0.7627
CKM constants no units $\pm\varepsilon$ $\pm\varepsilon^{rel}$ (%)
12-mix angle $\theta_{12}$ 0.22501 0.00068 0.3022
23-mix angle $\theta_{23}$ 0.04183 0.00079 1.8886
13-mix angle $\theta_{13}$ 0.003732 $9\times 10^{-5}$ 2.4116
CP-viol. phase $\delta$ 1.147 0.026 2.2668
Particle masses MeV $\pm\varepsilon$ $\pm\varepsilon^{rel}$ (%)
electron mass $m_{e}$ 0.510998 $10^{-6}$ 0.0002
muon mass $m_{\mu}$ 105.658 0.001 0.0009
$\tau$ mass $m_{\tau}$ 1776.93 0.09 0.0051
$u$-quark mass $m_{u}$ 2.16 0.07 3.2407
$d$-quark mass $m_{d}$ 4.70 0.07 1.4894
$s$-quark mass $m_{s}$ 93.5 0.8 0.8556
$c$-quark mass $m_{c}$ 1273.0 4.6 0.3614
$b$-quark mass $m_{b}$ 4183 7 0.1673
$t$-quark mass $m_{t}$ 172,560 310 0.1796
$Z$-boson mass $m_{Z}$ 91,188.0 2.0 0.0022
$W$-boson mass $m_{W}$ 80,369.2 13.3 0.0165
$H$-boson mass $m_{H}$ 125,200 110 0.0879
The input to the GP algorithm includes two dimensionless constants whose precision significantly surpasses that of the SM particle masses: the mathematical constant $\pi$ and the fine-structure constant $\alpha$ (or its inverse value $\alpha^{-1}$, which will be used in our study). Although the constant $\pi$ is not a parameter of the SM, it frequently appears in the laws of physics. To maintain consistency, we reduce their precision to $\pi=3.14159(1)$ and $\alpha^{-1}=137.036(1)$. Increasing the precision of $\pi$ and $\alpha^{-1}$ is not expected to affect the results involving particle masses, as the overall uncertainty will still be dominated by uncertainties in the mass values. However, using overly precise constants can introduce numerical instability and make it more difficult for GP to reproduce known relationships. These validation tests will be discussed later.
The main limitation of the inputs that contain a mix of constants with and without physical units is that only relations that pass dimensional analysis have physical meaning. There is no native support for dimensional analyses in the GP. Thus, the created GP output needs to be filtered out to accept only relations where physical units are the same on the left- and the right-hand sides of the relations. This results in significant inefficiency in the current algorithm’s generation of useful analytic relations.
The second challenge for this algorithm lies in the wide range of SM values, spanning several orders of magnitude—from the electron mass to the top quark mass, the latter being the heaviest among all SM particles. The uncertainties associated with the SM parameters vary significantly and must be taken into account when evaluating the robustness of any discovered relationships. If one variable has very small uncertainties, it offers little advantage if other input variables have large uncertainties, which will dominate the overall uncertainty on the output values. The least precise masses among our inputs are those of the light-flavor quarks ($u$, $d$, $s$), which were taken from the PDG using the $\overline{\text{MS}}$ scheme at a renormalization scale of 2 GeV. The most significant challenge is also when using the CKM mixing angles as inputs. Due to their relatively large uncertainties, there is a considerable risk of generating spurious equations.
It is important to note that high numerical precision does not necessarily translate to theoretical significance. For example, although the number $\pi$ can be calculated at an extraordinary number of decimal places, it is still just an irrational number, and the usage of all its digits is not essential for a physical theory, since digits beyond a certain level of precision have no observable effect on physical processes.
3.3 Inputs Without Physical Units
For the second set of GP inputs, we converted all masses shown in Table 1 to values without physical units. There are arguments [Duff_2014] suggesting that the laws of physics should be independent of the choice of units or measuring instruments. Dimensionless constants, in particular, are well-suited for probing the deep structure underlying physical laws.
@@ 3.3 Inputs Without Physical Units
One way to address the physical units issues is to normalize all particle masses by dividing them by another parameter that has a unit of mass. However, this approach introduces a degree of arbitrariness in the choice of the reference mass. A possible option for the divider is the Planck mass [ParticleDataGroup:2024cfk], $m_{\mathcal{P}}=1.22089(6)\times 10^{19}$ GeV. While theoretically appealing, this choice is impractical for our purposes, as the GP algorithm we employ cannot handle values smaller than $10^{-17}$.
@@
To resolve this, we normalize all SM masses by the mass of the $\rho(770)$ meson, $775.26\pm 0.23$ MeV [ParticleDataGroup:2024cfk]. At first glance, this might seem like an unconventional choice. One could also use the mass of any fundamental particle for normalization, such as the electron mass. However, that particle would then be excluded from our GP analysis. Our choice is not without justification for the GP algorithm. The $\rho(770)$ meson holds a special place in the context of quantum chromodynamics, plays a key role in hadronic structure, and even electromagnetic interactions. Its mass lies conveniently in the GeV range, which allows us to rescale the SM mass spectrum to a range that is well-suited for GP.
@@
For example, the mass of the electron, 0.51099895000(15) MeV, has an exceptionally small relative uncertainty of $3\times 10^{-8}\%$, while the Higgs boson mass, $125.20(11)$ GeV [ParticleDataGroup:2024cfk], has a much larger relative uncertainty of approximately 0.09%. This disparity in precision poses a challenge for numerical algorithms. However, scaling all masses by the $\rho(770)$ meson mass and propagating its uncertainty to the resulting ratios helps mitigate this problem. After this division, the values used for $\pi$ and $\alpha^{-1}$ remain 40 to 100 times more precise than the rescaled electron mass by the $\rho$-meson mass.
Table 2 presents the masses of the SM particles normalized by the mass of the $\rho$-meson. For consistency, we use the same symbol $m$ to denote the rescaled masses as for the original masses (with units). Thus, for the second set of inputs, we use the constants from Table 1 but where all masses are replaced by the values listed in Table 2.
Table 2: The masses [ParticleDataGroup:2024cfk] of the fundamental particles are rescaled to obtain dimensionless values, eliminating the mass units (MeV). For this purpose, all particle masses are divided by the mass $775.26\pm 0.23$ MeV of the $\rho$ meson to improve the value range and reduce the large variability of uncertainties, making them more uniform for GP. The traditional symbol $m$ was chosen for the rescaled masses to simplify the presentation of the results. The table presents both the absolute measured uncertainties ($\pm\varepsilon$) and the relative uncertainties $(\pm\varepsilon^{rel}$, in percent).
Constant Name Value $\pm\varepsilon$ $\pm\varepsilon^{rel}$ (%)
electron mass $m_{e}$ 0.0006591 2 × $\times\,10^{-7}$ 0.0303
muon mass $m_{\mu}$ 0.1363 4 × $\times\,10^{-5}$ 0.0293
$\tau$ mass $m_{\tau}$ 2.292 0.0007 0.0305
$u$-quark mass $m_{u}$ 0.002786 9 × $\times\,10^{-5}$ 3.2304
$d$-quark mass $m_{d}$ 0.006062 9 × $\times\,10^{-5}$ 1.4847
$s$-quark mass $m_{s}$ 0.1206 0.0010 0.8292
$c$-quark mass $m_{c}$ 1.642 0.006 0.3593
@@
We expect the number of valid relations reported by GP for inputs without units to be significantly larger than for inputs with physical units. This is primarily because the output and input variables are unitless. Therefore, all relations identified by GP will be considered valid. In addition, the inputs with the rescaled masses have significantly more uniform relative uncertainties.
It should be emphasized that the presence of the mass of $\rho(770)$ (or any other particle mass used for obtaining dimensionless SM constants) should be considered a handy method for deriving short relationships in GP. It is an auxiliary parameter. The final theory can be free of this constant after successful variable substitutions, which should be performed after obtaining the analytic expressions. As we discuss later, this means that the dimensionless input allows for the generation of analytically more complex and, at the same time, dimensionally consistent expressions. This will lead to a large functional coverage of the potential relations with relatively low CPU usage during GP production.
3.4 The GP Search Algorithm
The input parameters discussed above will be used for the following studies as described below.